# IFT-UAM/CSIC-25-108

### v1

@@ I Variation in chi-squared weights and model-selection criteria
$\mathrm{Tay2}][a^{2}][-]$ 0.70 0.0312 0.1451(8)
$\mathrm{Tay2}][a^{2}][\beta>3.40]$ 0.61 0.0159 0.1449(12)
$\mathrm{Tay2}][a^{2}][\beta>3.46]$ 0.45 0.0072 0.1451(14)
$\mathrm{Tay2}][a^{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.67 0.0206 0.1449(9)
$\mathrm{Tay2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.42 0.0049 0.1449(16)
$\mathrm{Tay2}][a^{2}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.33 0.0015 0.1449(10)
$\mathrm{Tay2}][a^{2}][m_{\pi}L>4.1]$ 0.84 0.0174 0.1453(9)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.70 0.0312 0.1451(9)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.60 0.0159 0.1449(13)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.47 0.0072 0.1451(14)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.67 0.0205 0.1449(9)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.43 0.0050 0.1449(16)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.33 0.0015 0.1449(10)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.84 0.0174 0.1453(10)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][-]$ 0.69 0.0272 0.1445(13)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.56 0.0118 0.1441(20)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.39 0.0040 0.1451(23)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.61 0.0120 0.1446(15)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.87 0.0157 0.1440(20)
$\mathrm{Tay4}][a^{2}][-]$ 0.72 0.0235 0.1446(9)
$\mathrm{Tay4}][a^{2}][\beta>3.40]$ 0.62 0.0101 0.1444(13)
$\mathrm{Tay4}][a^{2}][\beta>3.46]$ 0.45 0.0045 0.1447(15)
$\mathrm{Tay4}][a^{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.66 0.0111 0.1445(10)
$\mathrm{Tay4}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.36 0.0019 0.1446(17)
$\mathrm{Tay4}][a^{2}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.97 0.0053 0.1432(11)
$\mathrm{Tay4}][a^{2}][m_{\pi}L>4.1]$ 0.76 0.0080 0.1455(16)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][-]$ 0.67 0.0149 0.1444(12)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.53 0.0057 0.1441(20)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.38 0.0023 0.1455(23)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.58 0.0063 0.1447(15)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.92 0.0130 0.1439(20)
@@
$\chi\mathrm{SU(3)}][a^{2}][-]$ 0.84 0.0458 0.1444(4)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40]$ 0.80 0.0196 0.1442(6)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.46]$ 0.74 0.0032 0.1441(8)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.80 0.0081 0.1443(4)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.66 0.0010 0.1440(9)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.65 0.0002 0.1442(6)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}L>4.1]$ 0.77 0.0019 0.1445(5)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.84 0.0478 0.1445(4)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.79 0.0184 0.1443(7)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.75 0.0032 0.1441(9)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.80 0.0082 0.1443(5)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.67 0.0010 0.1440(9)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.64 0.0002 0.1443(6)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.76 0.0018 0.1446(5)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.89 0.0640 0.1440(5)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.83 0.0169 0.1439(7)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.72 0.0020 0.1438(9)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.81 0.0063 0.1440(5)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.62 0.0004 0.1438(9)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.57 0.0001 0.1440(6)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.85 0.0026 0.1440(6)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][[-]$ 0.86 0.0371 0.1446(5)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.80 0.0134 0.1445(7)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.74 0.0020 0.1444(9)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.82 0.0079 0.1445(5)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.63 0.0004 0.1442(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.56 0.0001 0.1443(7)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.80 0.0015 0.1449(7)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][-]$ 0.88 0.0457 0.1439(8)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.79 0.0081 0.1437(13)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.68 0.0010 0.1444(16)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.78 0.0035 0.1442(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.80 0.0013 0.1439(13)
@@
$\chi\mathrm{SU(2)}][a^{2}][-]$ 0.82 0.0276 0.1444(8)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40]$ 0.76 0.0104 0.1442(11)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.46]$ 0.71 0.0019 0.1442(11)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.78 0.0053 0.1445(8)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}L>4.1]$ 0.72 0.0011 0.1447(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.83 0.0288 0.1444(8)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.75 0.0098 0.1443(11)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.70 0.0018 0.1442(11)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\,\mathrm{MeV}]$ 0.78 0.0053 0.1446(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.71 0.0010 0.1447(9)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.87 0.0340 0.1440(8)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.80 0.0094 0.1439(11)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.69 0.0012 0.1439(12)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.82 0.0058 0.1442(9)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.82 0.0015 0.1442(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][-]$ 0.84 0.0229 0.1445(8)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.76 0.0059 0.1444(11)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.69 0.0013 0.1444(12)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.79 0.0036 0.1446(8)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.76 0.0009 0.1449(9)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][-]$ 0.86 0.0243 0.1438(11)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.76 0.0040 0.1436(16)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.61 0.0005 0.1443(20)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.80 0.0029 0.1438(13)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.82 0.0012 0.1433(17)

### v2

@@ I Variation in chi-squared weights and model-selection criteria
$\mathrm{Tay2}][a^{2}][-]$ 0.65 0.0377 0.1447(7)
$\mathrm{Tay2}][a^{2}][\beta>3.40]$ 0.57 0.0214 0.1445(9)
$\mathrm{Tay2}][a^{2}][\beta>3.46]$ 0.46 0.0079 0.1444(10)
$\mathrm{Tay2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.62 0.0198 0.1445(7)
$\mathrm{Tay2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.38 0.0050 0.1442(11)
$\mathrm{Tay2}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.36 0.0016 0.1447(8)
$\mathrm{Tay2}][a^{2}][m_{\pi}L>4.1]$ 0.79 0.0133 0.1447(7)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.66 0.0376 0.1447(7)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.57 0.0215 0.1445(9)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.46 0.0073 0.1444(10)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.61 0.0194 0.1444(7)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.38 0.0051 0.1442(11)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<350\;\rm{MeV}]$ 0.34 0.0016 0.1446(8)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.79 0.0138 0.1447(7)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][-]$ 0.68 0.0316 0.1438(10)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.54 0.0122 0.1436(14)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.36 0.0045 0.1446(20)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.54 0.0107 0.1442(13)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.82 0.0152 0.1432(15)
$\mathrm{Tay4}][a^{2}][-]$ 0.68 0.0205 0.1441(9)
$\mathrm{Tay4}][a^{2}][\beta>3.40]$ 0.55 0.0102 0.1440(11)
$\mathrm{Tay4}][a^{2}][\beta>3.46]$ 0.40 0.0036 0.1440(12)
$\mathrm{Tay4}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.59 0.0087 0.1440(9)
$\mathrm{Tay4}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.29 0.0018 0.1438(13)
$\mathrm{Tay4}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.99 0.0047 0.1430(11)
$\mathrm{Tay4}][a^{2}][m_{\pi}L>4.1]$ 0.71 0.0062 0.1451(13)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][-]$ 0.65 0.0170 0.1437(11)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.49 0.0054 0.1435(14)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.33 0.0023 0.1448(19)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.51 0.0051 0.1442(13)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.95 0.0122 0.1433(16)
@@
$\chi\mathrm{SU(3)}][a^{2}][-]$ 0.80 0.0484 0.1441(3)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40]$ 0.74 0.0149 0.1439(5)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.46]$ 0.74 0.0028 0.1436(7)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.81 0.0123 0.1439(3)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.71 0.0010 0.1434(8)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.69 0.0001 0.1441(4)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}L>4.1]$ 0.66 0.0012 0.1441(4)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.77 0.0423 0.1441(3)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.73 0.0124 0.1439(5)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.73 0.0027 0.1436(7)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.79 0.0113 0.1439(4)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.71 0.0011 0.1434(8)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<350\;\rm{MeV}]$ 0.70 0.0001 0.1442(4)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.64 0.0010 0.1441(4)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.89 0.0703 0.1437(4)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.83 0.0187 0.1434(6)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.75 0.0026 0.1433(7)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.83 0.0094 0.1437(4)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.67 0.0007 0.1433(8)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<350\;\rm{MeV}]$ 0.66 0.0001 0.1439(5)
$\chi\mathrm{SU(3)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.80 0.0017 0.1436(5)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][[-]$ 0.85 0.0404 0.1444(4)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.79 0.0108 0.1442(5)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.75 0.0022 0.1439(7)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.82 0.0103 0.1442(4)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.70 0.0004 0.1437(8)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.61 0.000 0.1442(5)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.75 0.0011 0.1446(5)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][-]$ 0.87 0.0523 0.1435(7)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.80 0.0139 0.1433(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.72 0.0008 0.1438(14)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.79 0.0048 0.1438(9)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.77 0.0009 0.1435(10)
@@
$\chi\mathrm{SU(2)}][a^{2}][-]$ 0.77 0.0255 0.1440(8)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40]$ 0.70 0.0080 0.1438(9)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.46]$ 0.72 0.0021 0.1437(10)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.78 0.0061 0.1440(8)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}L>4.1]$ 0.62 0.0007 0.1442(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.75 0.0218 0.1440(8)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.69 0.0067 0.1438(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.71 0.0019 0.1437(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.77 0.0057 0.1440(8)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.60 0.0005 0.1442(9)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][-]$ 0.88 0.0463 0.1437(8)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.81 0.0117 0.1435(10)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.74 0.0012 0.1434(10)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.82 0.0042 0.1438(8)
$\chi\mathrm{SU(2)}][a^{2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.80 0.0011 0.1439(9)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][-]$ 0.82 0.0198 0.1442(8)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.40]$ 0.76 0.0093 0.1441(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][\beta>3.46]$ 0.73 0.0013 0.1440(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.80 0.0038 0.1442(8)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][a^{2}][m_{\pi}L>4.1]$ 0.70 0.0005 0.1446(9)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][-]$ 0.86 0.0261 0.1434(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.78 0.0074 0.1431(13)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.68 0.0005 0.1436(18)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.79 0.0028 0.1435(12)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.80 0.0010 0.1429(14)